OpenStudy (anonymous):

Rationalize the denominator and simplify: 6/(square root of 5 - square root of 2)

OpenStudy (radar):

\[6/(\sqrt{5}-\sqrt{2})\]

OpenStudy (radar):

Multiply both the numerator and denominator by:\[\sqrt{5}+\sqrt{2}\]

OpenStudy (anonymous):

Do I distribute 6?

OpenStudy (anonymous):

Or, write it as 6(square root of 5 + square root of 2)?

OpenStudy (radar):